OpenStudy (anonymous):

Jim is three times as old as Allison. Two years ago, the sum of their age of 16. - Write a linear system

OpenStudy (zzr0ck3r):

3j=s (s-2)+(j-2) = 16?

OpenStudy (lgbasallote):

let allison's age = x jim's age = 3x two years ago their sum is 16 so jim's age 2 years ago = 3x - 2 allison's age 2 years ago = x - 2 so we'll have (3x - 2) + (x - 2) = 16 got it?

OpenStudy (zzr0ck3r):

" the sum of their age of 16."?

OpenStudy (anonymous):

oh okay, thank you !

OpenStudy (anonymous):

nope wait, where did you get 16 from ?

OpenStudy (lgbasallote):

sum of their age is 16

OpenStudy (anonymous):

omg sorry! nevermind

OpenStudy (lgbasallote):

hehe :)

OpenStudy (anonymous):

thank you!

OpenStudy (lgbasallote):

<tips hat>